5. Real versus nominal GDP Consider a simple economy that produces two goods: pencils and envelopes. The following table showFrom 2017 to 2018, nominal GDP C, and real GDP decreased The inflation rate in 2018 was increasedThe inflation rate in 2018 was -27% Why is real GDP a more accura 0.3% ure of an . Real GDP is not influen 27% rice chan effe
Answer #1

Real GDP is more accurate measure of economic growth and production than the nominal GDP because it is measured from the base price level .

So price is constant and taken from the base year .

For real GDP inflation factoris also taken.

Nominal GDP is calculated for the currently year always.

So the true answer is option A

hince, Nomiral Gap - Parreninger X qevererged (NGHI NADP 2016 - IX 125+ 1x 200 = $325 NGOP 2017 = P201A+22012 = 2x1707 4x230Now! Ramp 2017 : (2016) - 9 2018 [price will be of bare year = LXtot 1x230 = $400 Ramp 2018= P2016) X9 2018 = 12150 + 1x170 =27 Eppd2018 - (1280x100 3400 Now, from 2017-2018 Napp increased by: - - | Narp2018 - Nadprost 1x100 NGRP2017 = 1280-1260 ( 12